OpenStudy (andrea):

an investment of $2000 doubles every 12 yr. how much is it worth after 36 yr? after 60yr?

OpenStudy (anonymous):

$16,000 in 36 years and $64,000 in 60 years.

OpenStudy (anonymous):

2,000*2=4,000*2=16,000 36 years 2,000*2=4,000*2=16,000+2=32,000*2=64,000 60 years

OpenStudy (andrea):

Thanks